OpenStudy (anonymous):

Given that the pair of triangles is similar, find the length of the side labeled x. 1st triangle is 22.8 & x 2nd triangle is 5.7 & 9.7

Directrix (directrix):

22.8/5.7 = x/ 9.7 x = 38.8

OpenStudy (anonymous):

x = 22.8 * 9.7/ 5.7 = 38.8